The time needed for college students to complete a certain paper and pencil maze follows a Normal distribution, with a mean of 70 seconds and a standard deviation of 15 seconds. You wish to see if the mean time μ is changed by meditation, so you have a group of 9 college students meditate for 30 minutes and then complete the maze.Question 29 of 40> The time needed for college students to complete a certain paper and pencil maze follows a Normal distribu
Answer #1

Solution:

The test statistic z is given by

z = \frac{\bar x - \mu0}{\sigma/\sqrt{n}}

= (78.55 - 70) / (15/\sqrt{}9)

= 1.71

Now , observe that ,there is  \neq sign in H1. So , the test is two tailed.

For two tailed test :

p value = 2 * P(Z < -z)

= 2 * P(Z < -1.71)

= 2 * 0.0436

= 0.0872

Answer : between 0.05 and 0.1